1. Hero Manufacturing has 9 million shares of common stock outstanding. The current share price is...

1.

Hero Manufacturing has 9 million shares of common stock outstanding. The current share price is $75 and the book value per share is $6. The company also has two bond issues outstanding, both with semiannual coupons. The first bond issue has a face value $85 million and a coupon of 10 percent and sells for 96 percent of par. The second issue has a face value of $65 million and a coupon of 11 percent and sells for 109 percent of par. The first issue matures in 25 years, the second in 9 years.

a.

What are the company's capital structure weights on a book value basis? (Do not round intermediate calculations and round your answers to 4 decimal places, e.g., .1616.)

b. What are the company's capital structure weights on a market value basis? (Do not round intermediate calculations and round your answers to 4 decimal places, e.g., .1616.)

   

c.

Which are more relevant?

   

  • Market value weights

  • Book value weights

2.

Shanken Corp. issued a 20-year, 5.8 percent semiannual bond 4 years ago. The bond currently sells for 109 percent of its face value. The book value of the debt issue is $40 million. In addition, the company has a second debt issue on the market, a zero coupon bond with 11 years left to maturity; the book value of this issue is $40 million and the bonds sell for 53 percent of par. The company’s tax rate is 23 percent.

a.

What is the company's total book value of debt? (Do not round intermediate calculations and enter your answer in dollars, not millions of dollars, e.g., 1,234,567.)

b.

What is the company's total market value of debt? (Do not round intermediate calculations and enter your answer in dollars, not millions of dollars, e.g., 1,234,567.)

c.

What is your best estimate of the aftertax cost of debt? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)
Answer #1

1

a

BV of equity=Price of equity*number of shares outstanding
BV of equity=6*9000000
=54000000
BV of Bond1=Par value*bonds outstanding*%age of par
BV of Bond1=1000*85000*1
=85000000
BV of Bond2=Par value*bonds outstanding*%age of par
BV of Bond2=1000*65000*1
=65000000
BV of firm = BV of Equity + BV of Bond1+ BV of Bond 2
=54000000+85000000+65000000
=204000000
Weight of equity = BV of Equity/BV of firm
Weight of equity = 54000000/204000000
W(E)=0.2647
Weight of debt = BV of Bond/BV of firm
Weight of debt = 150000000/204000000
W(D)=0.7353

b

MV of equity=Price of equity*number of shares outstanding
MV of equity=75*9000000
=675000000
MV of Bond1=Par value*bonds outstanding*%age of par
MV of Bond1=1000*85000*0.96
=81600000
MV of Bond2=Par value*bonds outstanding*%age of par
MV of Bond2=1000*65000*1.09
=70850000
MV of firm = MV of Equity + MV of Bond1+ MV of Bond 2
=675000000+81600000+70850000
=827450000
Weight of equity = MV of Equity/MV of firm
Weight of equity = 675000000/827450000
W(E)=0.8158
Weight of debt = MV of Bond/MV of firm
Weight of debt = 152450000/827450000
W(D)=0.1842

c

MV weights are more relevant as they give picture of the current capital structure of the firm which impacts the cost of capital

Please ask remaining parts seperately, questions are unrelated